Agenda/Minutes of the VIS Meeting on 2016/11/11
2016/11/11 13:30 -
Participants: Shoda, Kasuya, Akutsu, Mark, Hirata, Okutomi, Aso, Fujii, Sato, Ishizaki, Tanaka, Ohishi, Enzo, Raffaele, Takahashi
- Installation: Asking a company the cost and schedule for transportation of Type-A items from the center area to the 2F of Y-end.
- Pre-isolator: Assembly work will be concentrated to the next week. The PI for SRs will be transported on 21st from Akeno. 5 ACCs were measured at the site.
- EQ stop, release sensor, dummy payload: Production is on going.
- BF damper: Open bidding was on 9th. A production company was decided.
- Maraging wire: Asking the company a quotation. The company is checking procurement of the material.
Type B (Mark)
- Type B Past Week Summary
- Mark and Hirata-san at Mitaka
- Gathered parts for cable making.
- Mark and Hirata-san at Kamioka:
- Replaced the floor panels that had been taken from the PR2 area (filling in the hole in the floor) and found/installed two new panels to act as bridges near the BS assembly frame.
- Installed two more LPCDs and the stepper motor driver in the PR2 rack.
- Ran cables from the rack to the second and third sat amps (for the IRM OSEMs).
- Added a second DIO card in the rack and connected it to the BIO inputs and outputs of the new LPCDs.
- Started making extension cables for RM-H2 and -H3 OSEMs but gave up due to complicated grounding arrangement.
- Washed four 40-cm D-Sub-9 ribbon cables to use as temporary extensions (maybe also for H1/H4).
- Added extension to the IM pico cable for the BF yaw pico.
- Made LVDT adapter cable.
- BS mirror box was shipped back.
- Mark worked on schedule.
- Mark and Hirata-san at Mitaka
- New flags, OB related parts, and the assembly frame were ordered. We are now asking the delivery time.
- Flag removing jigs
- The parts for the jigs are delivered at Mitaka. Some parts are passed to Fukushima-san for machining.
- TM OSEM flag
- We found the one of the TM OSEM flags are detached from the mirror when we opened the vacuum tank. Since TFs that Fujii-kun measured were not weird, it seems to be detached after the measurement. (Will be reported next week again.)
- The clump plates for IM may have to be re-manufactured. The current ones have pressed lines by wires.
- Put a report here
Watchdog system (Tanaka)
Schematic and PCB designJGW-E1605760-v2(PDF)
- Wearing a terminating resistor(for OSEM sensor deferential out) on both sides.
- And the value of terminating resistor is changed from 3k to 30k.
Design and preparation(investigation of procurement) for production.JGW-E1605791-v4(PDF)
- A strong incandescent light was put close to the plastic sheets of the clean booth, which is probably flammable.
- A temporary pillar in the PR clean booth near MCF came off when people moved the coach filters and hit it.
- Foreseen risks
- Type-B team will lift the suspension system from the BF to measure the weight. Mark will write up a procedure and circulate to the VIS people.
- Type-A team will continue the assembly of the PI. A crane work to combine the IP and TF is foreseen. Make sure to have at least two people doing the work.
- JPS/ASJ meetings
- VIS overall talk: Takahashi (JPS)
- TYpe-A: Okutomi (JPS)
- Type-B: Mark (JPS)
- Type-Bp: Shoda (JPS)
- Fujii may present his data analysis work at ASJ
- PR3 accident
- What happened
- When the suspension system was put back to the assembly frame, the position and height of the IM was not correct.
- During the adjustment process of the IM position, the mirror moved too much in roll.
- The mirror slept out of the wires and hit the mirror box (MB)
- MB and the recoil mass (RM) moved on the rail by the shock creating a gap between them.
- The mirror fell in the gap between the MB and RM.
- What was wrong
- The assembly jigs were designed only with the assembly procedure in mind. Disassembly was not taken into account at the time.
- Insufficient adjustment mechanism of the IM position.
- Long rod from the mirror box interfered with the work
- No slit on one wire breaker
- Base for the RM is not fixed. RM was not fixed to the base.
- MB base and the RM base were not connected together.
- Did not review the disassembly procedure beforehand.
- Adjustment of the mirror position should have been done while the mirror is inside the RM.
- Adjustment of the mirror position took 2 hours. People got tired.
- There was only one expert in the team. Should have had at least two.
- The allocated time for the PR3 disassembly was too short.
- Communicating in English while doing a complicated assembly work was difficult.
- Needed to take more rest.
- We will discuss on how to solve the technical problems as well as increase the overall safety of the VIS works in the future at the next meeting.
- Shoda will write a report on this accident: What happened, What was wrong, and what are we going to do.
- What happened
Site work plan for the next week
- Wed: BS
- Thu: BS
- Fri: BS
- Travel (Week of 11/14):
- Mark: 16-18
- Hirata: 16-18
- Enzo: 16-18
- Travel (Week of 11/21):
- Takahashi: 21 Akeno
On 2016/11/18(Fri) in Mitaka
VIS F2F: Nov. 18, Dec. 2